Ace Industrial Machines issued 190,000 zero coupon bonds 7 years ago. The bonds originally had 30 years to maturity with a yield to maturity of 6.6 percent. Interest rates have recently decreased, and the bonds now have a yield to maturity of 5.7 percent. The bonds have a par value of $2,000. If the company has a $88.2 million market value of equity, what weight should it use for debt when calculating the cost of capital? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., .1616.)
Answer #1

Weight of bond = Revised market value of the bond / Total market value of capital

Revised market value of zero coupon bond =

Par value / (1 + R)^N

par value = 2000

R = 5.7%

N = 30 -7 = 23

Revised market value = 2000 / (1 +5.70%)^23

= 558.86

Value of 19000 bonds = 190000 * 558.86 / 1000000 = $106.1834 million

Total market capital = 88.2 + 106.1834 = 194.3834 million

Weight of debt = 106.1834 / 194.3834

= 54.6258%